OpenStudy (anonymous):

How long would it take to double an investment of $1000 if it is invested at 3.5%, compounded continuously?

OpenStudy (anonymous):

Solve 1000*1.035^n = 2000 ... which simplifies to 1.035^n = 2

OpenStudy (anonymous):

You know how to solve that? Use log's.

OpenStudy (anonymous):

what formula of logs would be used?

OpenStudy (anonymous):

Take the log of both sides, first to get log(1.035^n) = log(2) Then use the fact that log(a^b) = b*log(a)... n*log(1.035) = log(2) n = log(2) / log(1.035) =_________